You want to create a portfolio equally as risky as the market, and you have $500,000 to invest. Information about the possible investments is given below:

  Asset Investment Beta
  Stock A $ 144,000       .89      
  Stock B $ 136,000       1.34      
  Stock C 1.49      

  Risk-free asset

1. How much will you invest in Stock C? (Do not round intermediate calculations. Round your answer to 2 decimal places (e.g., 32.16).

2. How much will you invest in the risk free asset?
Answer #1

(144*0.89 + 136*1.34 + x*1.49) = 500*1

X or the amount in C = 127,248.32

amount in risk free asset = 500,000 - 144,000 - 136,000 - 127,248.32 = 92,751.68